Conduct Hypothesis Testing Using the Traditional Method

The Mathematics Department in a certain university is conducting a study to determine how long it takes its graduates to find a job. A sample of 80 graduates was surveyed and it was found out that the average time it has taken a graduate to find a new job is 2.5 months, with a standard deviation of 2 months. Is there a sufficient evidence to conclude that the graduate of this department takes on the average more than two months to find a job? Use the 90% confidence level. 




1
Expert's answer
2022-03-20T06:44:19-0400

"\mu=2.5 \plusmn z_{\alpha/2}\frac{\sigma}{\sqrt{n}}=2.5 \plusmn 1.645 \frac{2}{\sqrt{80}}=2.5 \plusmn{0.37}"

Answer: Yes it takes from 2.13 to 2.87 month